OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (1, 2) and (2, 5)?

OpenStudy (anonymous):

Well first find the slope, using the following formula: \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where your coordinates represent \[(x_1,y_1)~~\text{and}~~(x_2,y_2)\] Once you manage to get that, plug in m and your first coordinates in the point - slope form: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

ok jus a sec

OpenStudy (anonymous):

so do i leave y blank n just plug in y1?

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

would y1 be the numerator of the slope

OpenStudy (anonymous):

i mean midpoint?

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

\[(1,2) \implies (x_1,y_1)\]

OpenStudy (anonymous):

You can use (x_2,y_2) as well if you wanted to instead it doesn't matter